Oliver Inc. acquired the following assets in January 2017:

Equipment: estimated useful life, 5 years; residual value, $15,000 $470,000
Building: estimated useful life, 30 years; no residual value $720,000


The equipment was depreciated using the double-declining-balance method for the first three years for financial reporting purposes. In 2020, the company decided to change the method of calculating depreciation for the equipment to the straight-line method, because of a change in the pattern of benefits received (but no change was made in the estimated useful life or residual value). It was also decided to change the building’s total estimated useful life from 30 years to 40 years, with no change in the estimated residual value. The building is depreciated using the straight-line method.

(a) Prepare the journal entry to record depreciation expense for the equipment in 2020. (Ignore tax effects.)

(b) Prepare the adjusting entries.
Answer #1
Date Account Description Debit($) Credit($)
Dec-31 Depreciation expense - Equipment 43,260.00
Accumulated Depreciation - Equipment 43,260.00
Dec-31 Depreciation expense - Buildings 17,513.51
Accumulated Depreciation - Buildings 17,513.51

* All the journal entries and workings are done presuming that the change in depreciation basis happened on Jan 1, 2022.

Notes

1) BUILDINGS

Cost of Building = $720000

Estimated Useful life = 30

(Original) Annual Depreciation under Straight Line Method = $720000/30 = $24000

Accumulated Depreciation for 3 years (2019-2021) = 24000*3 = $72000

When estimated life is changed to 40 years, 3 years have already elapsed, the remaining value of the asset will be depreciated over 37 years.

Revised Annual Depreciation = (720000-72000)/37 = $17513.51

2) EQUIPMENT

Cost of Equipment = $470000

Residual Value = $15000

Estimated useful life = 5 years

Depreciation schedule under Double declining method

Year Book Value Depreciation Depreciation Accumulated Book Value Expense Percent Year Start Depreciation Year End 2019 $470,0

Book value at the beginning of 2022 = $101520

Depreciation under Straight Line method = (101520-15000)/2 = $43260

Please comment in case of any issue and I will be happy to help.